Project

General

Profile

Actions

Ops/Comm Scheduling meeting : 21/3/2017

Location : CR03 in R1 at 3:30 pm

Phone access : New phonebridge number
UK: 08082380274
US: 631.621.5253
Access code: 8161978083
Leader code: 2735

  • Actions from previous meeting
  • DC update
  • LH2 update
  • List of Hall activities in the shutdown (Colin)
  • List of issues arising from previous user cycle (may overlap with Colin's list) :
  • Recurring list (may overlap with Colin's list)
  • Start-up plans for next user cycle
    • ISIS cycle : May 2 - June 2
    • Propose first week (May 2 - 5) for full system startup, debugging, beam tuning etc
      • Will have 3 shifts available for monitoring during this period
    • Data-taking on a 24 hour basis to start on May 6 through to June 2

AOB:

Minutes

  • Actions from previous meeting
    • JBoehm to get SS operating info from PHanlet (done)
    • JBoehm to get/impose consensus from SC Magnets group on running of "Central CC GUI" (done)
    • JBoehm arrange tests of polarity effect on SSD leaks (done, no effect on leaks)
    • Hall Manager when removing FC to make a crude check that LiH hasn't moved. (done - no obvious movement)
    • Hall Manager: water system change-over procedure (stands)
    • Status of miceraid5 disk cleanup (stands)
    • DC to ensure Health Physics have checked temporary neutron monitor (n/a)
    • SBoyd - ensure list of issues arising from current running is being collated (done)
    • CWhyte - draw up the shutdown task list (done)
    • SBoyd - ensure MICE-wide discussion of MOM C&M panel (stands)
    • PFranchini - Get all shifters to go through the CE checklist during Control Room refresher training (stands, N.B. also change to neutron monitor)
    • MUchida - investigate moving Tracker waveguides around to cure "hole" (stands)
    • AKurup - send water pump GUI password to Henry (done)
  • DC update
    • LH2 update
      Absorber vessel is now in AFC1 in the Hall
      Found and fixed a leak in a brazed pipe.
      Now leak checking again, especially the Indium seals
      Phil Warburton now working on the controls
      The UPS was tested last week and is OK
      The Hydrogen detection system was tested last week and is OK
      Need a mesh to catch wayward MLI if something fails
      It will take 5-6 days to fill with H2, experts will NOT be present 24/7 so will need some level of shifter cover. VBayliss to confirm with Steve Boyd.
    • Joe, John Cobb and Victoria looked at diffuser yesterday: were able to withdraw assembly sufficiently to access iris 4.
      Some of the screws holding tungsten to carrier had come loose; they have been Loctite'd back in.
      Also noticed some magnetisation of parts of the tungsten petals.
      The Diffuser has been tested without field; decision on whether or not to test with field will be made when the run-up schedule becomes clear.
    • FC power supply is away for repairs, at least a week. The second supply has been returned from DL.
    • John Govans is replacing the permanent pipework to the compressors in the Trench
    • The CMS-ION neutron monitor has been returned and the temporary WENDI taken away. Remote readout not yet working.
    • Tracker cryostats 1 and 2 are being warmed up for the cold head service - this will be 24hrs late due to cable issue.
      ISIS crew could not turn off cryostat 2 as requested: it ignored the front panel rocker switch. This is a feature of that specific remote control cable and was probably the cause of the overheating incidents.
  • List of Hall activities in the shutdown (Colin)
    Updates are on the page. A number of the items (labelled "???") are unclear.
    • Network errors messages – in progress
      Need to specify a port number for all CArepeaters - thus need to reboot all IOC-type machines. Need to find good time for miceecserv reboot.
      Possibly related to the Watchdog failure, which hasn't re-appeared.
    • We have also noticed a lot of unknown devices believed to be server IPMI cards broadcasting DHCP requests across the network
  • List of issues arising from previous user cycle (may overlap with Colin's list) :
    • Get a spare EMR board (stands)
    • Compressor 2 SSU communication problems (Ajit is aware; possibly a comms cable)
    • Communication dropouts in the FCD helium level monitor (stands)
    • TKU compressor 2 needs servicing/ spares (stands)
    • TKD may need to swap a waveguide to improve dead channel placement (stands)
    • TKU butterfly helium leak tightening (MU, stands)
    • Status of target1 backup machine in R78 (EO/JL, stands)
    • need to emphasise procedures to gain swipe access in the vac area for new shifters (Steve Boyd)(stands)
    • need a spare PCI board for the tracker ioc pc backup machine (stands)
    • MLCR network isolation test still needs to be done (PF, stands)
  • Recurring list (may overlap with Colin's list)
    • #1878, #1891 Trench water pump control - pressure appears stable; shifter lockout in progress
    • #1881 Water system PVs not being archived - Services/Environment/Computing spreadsheet (stands)
    • #1880 Alarm Handler not reacting to conventional magnet interlock (stands)
    • #1903, #1881 micestore disk full, also needs to be in Nagios warning checks in place but need testing
    • #1881 File Compactor failure - shifters reported only the "previous error" message; should be in Alarm Handler. Move File Compactor output to top of RC screen? Also waiting on Services/Environment/Computing spreadsheet (stands)
    • VME crates in Hall overheating; crate monitoring looking at wrong crate Services/Environment/Computing spreadsheet, GUI needs fixing (stands)
    • Is MOM DC Configuration Checklist Elog tab up-to-date? - No, but it should be
    • Neutron monitor & restart docs - MICE to take the CMS ION to Poole to be fixed ourselves. MLCR needs to know how to restart remote viewing of the temporary monitor.
  • Start-up plans for next user cycle
    Shifts will be available via CHEESE in next couple of days
  • AOB
    • Henry will reboot micestore tomorrow to pick up the new kernel - client machines may need their NFS mounts restarting.

Actions

  • Hall Manager: water system change-over procedure (stands)
  • Status of miceraid5 disk cleanup (stands)
  • SBoyd - ensure MICE-wide discussion of MOM C&M panel (stands)
  • PFranchini - Get all shifters to go through the CE checklist during Control Room refresher training (stands, N.B. also change to neutron monitor)
  • PFranchini - Look at MOM-DC Configuration eLog tab and update if possible
  • MUchida - investigate moving Tracker waveguides around to cure "hole" (stands)
  • MCourthold - show JBoehm how to use the Picologger for quench data
  • JBoehm - get the West Wall compressors correctly labelled, as various have been re-purposed
  • VBayliss - confirm shifter cover needed for LH2 startup with Steve Boyd
  • DC - neutron monitor remote readout

Updated by Nebrensky, Henry over 6 years ago · 14 revisions